OpenStudy (anonymous):

diameter of ferris wheel is 80 ft.

OpenStudy (anonymous):

a) if the ferris wheel made one revolution every 45 second , find the linear velocity of a person riding in the ferris wheel b) suppose the linear velocity of person riding in the ferris wheel is 8 feet per second .What is the time for one revolution of the ferris wheel ?

OpenStudy (anonymous):

(a) Distance traveled in one revolution (circumference) = \[2\pi r = 2\pi \times 80ft = 502.65ft\] Time elapsed in one revolution = 45 seconds. \[\frac{502.65ft}{45s} = 11.17ft/s\] (b) \[\frac{502.65ft}{8\frac{ft}{s}} = 62.83s\]

OpenStudy (anonymous):

Oh wait, the diameter is 80 ft, I was using r=80ft

OpenStudy (anonymous):

(a) 5.585ft/s (b) 31.416s

OpenStudy (anonymous):

I think you right from begining d= diameter * pi

OpenStudy (anonymous):

=80*pi

OpenStudy (anonymous):

Well I was doing 2*pi*80 because I thought 80 was the radius, when it only should have been pi*80

OpenStudy (anonymous):

80*pi=251.33?

OpenStudy (anonymous):

I don't get b ?